# Follow the Flow: On Information Flow Across Textual Tokens in Text-to-Image Models

### v1

@@ 4.2 Are all prompt tokens necessary?
Our results show that removing non-representative tokens generally does not harm the generation (see 3, left-hand side). Of the cases where the original generation was aligned with the prompt, in 98% of the cases, the generated image still aligns with the prompt, suggesting that these tokens are largely redundant. Moreover, in cases where the original image failed to align with the prompt, we observe a 21% reduction in error rate (see 3, right-hand side). We attribute this improvement to the model being guided more by the representative tokens after removing the non-representative ones. See 2 in A.3 for full results.
Can we identify redundant tokens without generating their images?
Identifying which tokens are representative can improve image generation by removing redundant ones as shown in 3 (right). Moreover, many existing T2I methods operate at the token level, typically using all of a lexical item’s tokens to represent the item. (Chefer et al., 2023a; Dahary et al., 2024). Refining these methods to operate only on representative tokens could lead to improvements without requiring any additional changes. However, generating an image for each token is computationally expensive. To this end, we propose two ways to estimate token importance without image generation. First, we calculate the edit distance between a lexical item (e.g., “giraffe”, which is tokenized as “gir”, “a” and “ffe”, and a given token (e.g., ‘a’). We find that higher edit distance indicates a lower likelihood that the token is representative (Pearson correlation of $-0.44$). While this correlation is relatively strong, it is not accurate enough for practical use. Instead, we train a $k$-NN classifier ($k{=}5$, Euclidean distance) to predict token redundancy from encoded representations. We extract 6,966 tokens (4,864 unique lexical items) from our dataset, each labeled for whether the VLM labeled this token as representative. We find that this simple classifier achieves a precision of 92% in predicting token redundancy, highlighting its potential as an efficient tool for removing non-representative tokens while preserving the essential ones. Further details about the classifier can be found in A.1.
5 Inter-item information flow
User prompts are typically composed of multiple lexical items. Here we investigate the influence of lexical items on one another. For example, in the sentence “a pool next to a table”, is the representation of the item “pool” affected by the presence of the item “table”? More generally, given a prompt, made up of several lexical items, we aim to measure whether the presence in the prompt of one item affects the representation of another. To this end, we conduct the following experiment: given a prompt, we isolate each lexical item one at a time. First, we encode the entire prompt. Next, for each item, we patch the representations of all other tokens using our method Section 3.1, and generate an image based on the modified sequence. Specifically, we define a subset $S\subset\{1,\ldots,N\}$ containing the token indices of the selected lexical item, construct the patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$, and use it to guide the diffusion model. This process produces one image per lexical item in the prompt—each encoded in the full context of the prompt but with all other token representations masked. These images allow us to assess what information is encoded in each item’s representation. We then generate an image for the uncontextualized version of each lexical item—encoding only that item in isolation, without the rest of the prompt. To assess whether a lexical item $y$ influences another item $x$, we use Qwen2-VL to check if $y$ is present in the image generated from $x$, both with and without context. We then compare the Qwen2-VL answers between the contextualized and uncontextualized versions. Refer to A.1 for more details regarding Qwen2-VL and the prompts we use. An influence is considered True if $x$ did not exhibit $y$ in the uncontextualized representation but did in the contextualized one. Confirming that $y$ does not appear in the isolated representation of $x$ is crucial, as we want to ensure that the presence of $y$ in $x$ is due to information flow that occurs during textual encoding, and not due to some underlying correlation between the items. This approach allows us to quantify the degree to which lexical items affect one another within a prompt. See 4 for a demonstration of this framework.
@@ 5 Inter-item information flow
Our results show that in 89% of the cases, lexical items are not influenced by other items in the prompt. In the remaining 11%, one lexical item incorporates information about another item that appears in the same prompt; see 5 for examples. Interestingly, we find that when such influence occurs, it often involves items that frequently co-occur in natural language, and are therefore likely to appear together in the training data. For example, the item “pool” often shifts its representation from a swimming pool to a pool table when “table” is present in the context. This phenomenon occurs even when “table” is not used as a pool table, as in the sentence “A pool next to a table” (see 4). To measure the extent of information flow between items that should not influence each other (as in the “pool” and “table” example), we divide the identified cases of flow between items into two categories. The first includes cases where the items are semantically linked, such as “black” and “bear” in the prompt “a black bear”. The second includes cases where the items are not semantically linked, such as “black” and “bear” in the sentence “a green bear by the black tree”. To achieve this, we use an LLM as a judge (details in A.1). Our analysis reveals that 31% out of the influenced instances exhibit unintentional information flow.
Application.
Based on our findings, we present a method for mitigating semantic leakage. Our approach first encodes the prompt as usual, then separately encodes the item suspected of leakage. Finally, we replace the representation of the contextualized, leaked item in the original prompt’s encoding with its uncontextualized representation, and condition the diffusion process on this mixed representation (see 6). To evaluate the effectiveness of our method, we adapt prompts from Rassin et al. (2022) as these prompts often induce semantic leakage. For example, given the prompt “a bat is flying over a baseball stadium”, models tend to generate a flying baseball bat rather than the animal, due to the strong association between “bat” and “baseball stadium”. Since the dataset contains only about 30 examples, we use GPT-4o to expand it to 110 prompts. See A.2 for details on the augmentation process, including the full list of prompts. Out of the 110 prompts, 104 exhibit semantic leakage (as determined by manual human evaluation), where the image contains a concept not intended by the prompt. We then select the lexical items in each prompt that we suspect of having leaked, and apply our method on them. Remarkably, we achieve an 85% improvement on this set. See 8 in A.3 for additional qualitative examples.
@@
Figure 6: Removing semantic leakage by replacing the contextually leaked concept representation. (1) Regular generation produces an image showing a crosswalk to the right of a bus station. (2) Generation from the prompt “standing zebra”, without any context, results in the correct interpretation of the zebra as an animal. (3) Generation using the original prompt, but with the leaked concept “standing zebra” replaced by it’s uncontextualized representation, yields a correct image.
6 Conclusions
In this study, we examined the information flow within lexical items and between different lexical items during the textual encoding of a prompt in T2I models. We introduced an intervention to evaluate how information is distributed across various token representations. For in-item information flow, our analysis revealed that information within a concept is not uniformly distributed; instead, it is primarily concentrated in a few specific tokens. For inter-item information flow, we found that information typically does not flow between different items. When it does, the flow is often unintended and may result in semantic leakage. Our experiments suggest that semantic leakage originates within the text encoder itself, challenging the common assumption that each token primarily captures its intended meaning. We showed that semantic leakage phenomenon can be largely mitigated by simply encoding the leaked item separately and patching it into the prompt representation before the diffusion. While prior work has focused on modifying cross-attention to address challenges in text-to-image models, our results suggest that these methods should also account for information flow between tokens that occurs earlier, during the textual encoding.
7 Acknowledgments
This research was supported in part by the Israel Science Foundation (grant No. 448/20), NSF-BSF grant 2020793, an Azrieli Foundation Early Career Faculty Fellowship, an AI Alignment grant from Open Philanthropy, the European Union (ERC, Control-LM, 101165402), and by an Academic Gift by NVIDIA. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.1 Technical details
Lexical item classification
@@
Redundant token classification
We propose a probing classifier to predict whether a token is redundant (i.e., non-representative of its lexical item) using solely its encoded representation, without the need to generate an image. For this purpose, we extracted the 6,966 tokens corresponding to 4,864 unique lexical items from our dataset. Each token was annotated with a binary label indicating whether it represents the lexical item by the VLM (see A.1 for more details). We split the data into training and validation sets using an 80-20 ratio. A $k$-nearest neighbors (k-NN) classifier with $k=5$ and Euclidean distance as the similarity measure was then applied to predict token redundancy directly from the encoded representations. The results on the evaluation set are presented in 1. The high precision indicates that, in practical settings, one can remove tokens predicted as redundant with a high degree of certainty that representative tokens will not be inadvertently discarded.
@@
Evaluation visual generations.
@@
”In Yes, No and maybe. Does every image match one of those descriptions: (description string)? Answer Yes if all images match or relate to at least one description, Maybe if only some match, otherwise No.”
@@
Evaluating generated textual descriptions.
We employ gpt4o (OpenAI et al., 2024) to evaluate the textual interpretations produced by Patchscopes. We use the following prompt:
@@
Classifying the relations between items.
We enhance our leakage validation by distinguishing between cases where two lexical items exhibiting semantic leakage are perceptually bound together—for example, “old” and “man” in the prompt “a portrait of an old man”—and cases where they are not as “cone hat” and “eating” in the prompt “A person searing a cone hat is eating” (see 5). To achieve this, we use a large language model (LLM) as a judge. Specifically, we employ the following prompt:
@@ A.2 Data
DrawBench (Saharia et al., 2022): We include all categories except for “misspelling”, “rare words”, and “text”. Overall we end up with 134 prompts from DrawBench. Parti Prompts (Yu et al., 2022a): We include all categories except “Style & Format”, “Writing & Symbols”, and “Arts”. Overall we end up with 923 prompts from Parti Prompts. In total, we end up with 1,056 prompts.
Extended dataset of leakage prompts
Our augmentation process incorporates two components. First, we generate variations of existing prompts from Rassin et al. (2022) (e.g., modifying ’a gentleman with a bow in the forest’ to ’a man wearing a bow in the jungle’). Second, we introduce novel prompts with potential semantic leakage. For these prompts, we applied a one-lexical item change test by generating an image from a similar prompt that substitutes the affected or leaked item with an alternative term (e.g., replacing ’bishops’ with ’cardinals’ or ’checkers’ in ’chess in ’2 bishops playing chess’). This test ensures that minimal lexical modifications do not alter the intended semantic meaning while producing a different image due to semantic leakage from another item in the prompt (see the first two columns in 8 for few visual examples) . Together, these methods enrich the dataset and provide a robust framework for analyzing semantic leakage. The full list of prompts is available in the following anonymous Git repository: https://anonymous.4open.science/r/TokenRole-3E19.
A.3 Additional results
Figure 7: Distribution of representative vs. non-representative tokens across different lexical item lengths.
Method Success Rate
Regular Generation 84.1%
After Removing Redundant Tokens 87.6%
Table 2: Image-text alignment success rate (evaluated with Qwen-VL): comparison between standard generation and generation with redundant tokens removed
@@ A.3 Additional results
A.4 Catastrophic negligence
Catastrophic negligence refers to cases where a concept or entity is completely absent from the generated image. Chefer et al. (2023a) attempted to address this issue by increasing attention to the neglected item, implicitly assuming that the model understands the concept. We find that in 7% of the cases, the item is not generated from any of its constituent tokens and is absent from the full prompt’s image. To better understand this phenomenon, we leverage Patchscopes (Ghandeharioun et al., 2024), a tool that interprets intermediate token representations by generating textual explanations using the decoder in encoder-decoder architectures. Using Patchscopes, we observe that in approximately 67% of these cases the tokens are explained coherently, indicating that the item is adequately represented in the text encoder. For example, when we patch the encoded representation of the lexical item tuba into the prompt “describe X”, Patchscopes describes it as “a musical instrument played by blowing into the mouthpiece.” However, when we try to generate an image from its representation, the model fails (see 9), suggesting that the diffusion model may have encountered too few examples of the intended entity during training. Additionally, we identify two abstract categories that consistently result in failure in image generation, while Patchscopes accurately describes the concept: negation and numbers. For instance, the lexical item “without” failed to generate the an image without the item in 100% of 100 observed instances. See  9.
Figure 9: Cases where intervention-based interpretations fail, while Patchscopes accurately describe token meaning. Each column corresponds to a different prompt. The first row shows images generated from the contextual bolded token. The second row presents the same token explained by Patchscopes, and the third row shows the final image generated from the full prompt. Left: The image generated from the token “tub” (from “tuba”) depicts a lamp, whereas Patchscopes correctly identifies it as a musical instrument. Middle: The concept “without” fails to yield an image lacking a refrigerator, although Patchscopes correctly describes its semantics. Right: The concept “7” leads to an incorrect number of entities in the image, while Patchscopes correctly identifies the quantity.
A.5 Additional models
A.5.1 FLUX-Dev
In addition to our primary experiments with FLUX, we repeated all analyses using the Flux-dev variant. The redundant versus representative token experiments yielded similar trends, with 55% of tokens identified as representative and 45% as non-representative—values closely matching those observed with FLUX. Likewise, our inter-item flow experiments confirmed that information flow occurred in 11% of cases (and 3.1% miss intended leakage), reinforcing the overall patterns reported in the main text. Notably, while the aggregate trends are consistent across models, the specific lexical items resolved can differ between FLUX-schnell and Flux-dev, indicating potential a slightly different inner-lexicon. These findings underscore the robustness of our approach while highlighting model-dependent nuances in token representation and information flow dynamics.
A.5.2 SDXL-Turbo

### v2

@@ 3.2 Tokens: Are They All Necessary?
Our results (Table˜1) show that removing non-representative tokens generally does not harm generation (see Fig.˜4, left-hand side). When the original generation is aligned with the prompt, the generated image after non-representative token removal remains aligned in 98% of cases, suggesting these tokens are largely redundant. Surprisingly, in cases where the original image fails to align with the prompt, we observe a 21% improvement in alignment after removing non-representative tokens (see Fig.˜4, right-hand side). We attribute this improvement to the model relying exclusively on the remaining representative tokens, which encode the correct semantics of the item. [^2] We measure improvement rates with Qwen2-VL before and after masking non-representative tokens. See Section A.1.3 for further details.
Figure 5: Comparing Patchscopes to image generation with token-level patching in cases of item negligence. We assess whether a token’s concept is preserved by comparing images generated from its contextual representation, Patchscopes’ textual interpretation, and the full prompt image. Left: “tub” (from “tuba”) is correctly described by Patchscopes but fails to ground visually. Right: “paw” (from “pawn”) has missing details by both interpretations, suggesting a gap in the encoder’s conceptual knowledge.
3.3 Item Negligence
So far, we have explored scenarios where an item lacks a representative token, often leading to its omission from the final generated output, i.e., item negligence (Chang et al., 2024b; Chefer et al., 2023a). We next investigate why this omission occurs. We examine two potential causes: either the text encoder poorly represents the concept; or the concept is well-encoded, but the diffusion model isn’t familiar with its visual appearance.
To study an item’s encoding without relying on the diffusion model, we use Patchscopes (Ghandeharioun et al., 2024), which decodes token representations into their natural language descriptions, using the full encoder-decoder architecture of T5 for text generation. Specifically, we patch an item’s encoding into the template “describe <item>” and evaluate whether the output indeed describes the item using GPT-4o (OpenAI et al. 2024; see Section˜A.1.4 for prompt details). We apply this process to all neglected lexical items found in Section˜3.1, and examine whether the encoder is familiar with these items.
We observe that in 67% of all instances of negligence, Patchscopes returns correct descriptions (for example, as demonstrated in Fig.˜5, “tuba” yields “a musical instrument played by blowing into the mouthpiece”). This suggests that, in such cases, item negligence is caused by a failure in the diffusion model. In the remaining 33%, the Patchscopes outputs are incorrect, indicating a gap in the encoder’s semantic understanding or knowledge.
This points to two distinct sources of item-level negligence: Representational gaps within the text encoder itself—consistent with findings that augmenting text-side item descriptions improves factual generation (Huang et al., 2025)—and visual grounding failures despite sufficient text encodings.
Category Count Percentage
# pairs 15,950 100.00%
No Information Flow 14,251 89.35%
Information Flow 1,699 10.65%
   – Source before Reference 0835 49.15%
   – Reference before Source 0823 48.44%
Table 2: Distribution of information flow between lexical item pairs. Subcategories under “With Information Flow” indicate the order of source and reference.
4 Cross-Item Interactions
In Section˜3, we demonstrated that a single token can effectively encapsulate the semantics of an entire lexical item. This raises a broader question: What defines the segmentation boundaries of a lexical item? Can its tokens further encode adjacent items, or even the general surrounding context?
To this end, we isolate each lexical item in a prompt and assess whether it encodes information about other items in the prompt. For each item, we generate images from its contextualized representation (encoded within the full prompt), and its uncontextualized representation (using the item’s text by itself as a prompt). We then use Qwen2-VL to evaluate whether any other item from the prompt appears in the contextualized image, but not in the uncontextualized one—indicating information flow introduced by the surrounding context. See Section˜A.4 for implementation details.
Our results (Table˜2) show that in 89% of the cases, lexical items do not encode information from other items in the prompt. In the remaining 11%, one lexical item incorporates information about another item (see Fig.˜6 for qualitative examples).Interestingly, information flow can emerge between items with no direct syntactic relation in the prompt.
To better understand the nature of information flow, we analyzed the cases where lexical items appeared to influence one another. We categorize these prompt into 2 groups, based on whether this influence aligns with the prompt’s syntactic structure or not. The first category involves items with a direct syntactic relationship, such as the adjective “black” modifying the noun “bear” in the prompt “a black bear”. The second category includes items that are syntactically unrelated, such as the adjective “red” influencing the noun “building” in “a red car next to a building”. Using GPT-4o to classify between these cases (see Appendix A.1.5), we find that 31% of these information flow instances occur between syntactically unrelated items.
@@ 4 Cross-Item Interactions
Wrong information flow leads to misinterpretation.
Our analysis reveals a recurring theme of semantic influence between items within a prompt, a pattern not prominent in most of the prompts. In these cases, the interpretation of one item is skewed by another semantically related, but syntactically distant, item. For instance, in the prompt “A standing zebra to the right of a city bus station”, the strong semantic connection between “bus station” and “zebra” leads the model to generate a “zebra crossing” instead of the animal (see Fig.˜7). We refer to this outcome, where contextual association overrides the prompt-intended meaning, as an incorrect item resolution.
We note that such failures typically occur in cases of polysemous words. For example, the item bats is interpreted as wooden baseball bats in the prompt “bats fly around a baseball stadium”, whereas the intended meaning is obviously the flying animal. In such cases, the context overrides the correct sense of the item, leading to inaccurate generation. We hypothesize that these cases of incorrect item resolution are caused by information flow from the context to the misinterpreted lexical item. To validate this, We use the dataset from Rassin et al. (2022), which includes prompts known to induce semantic leakage due to implicit lexical associations (e.g., “bat” and “baseball stadium”). Unlike standard T2I benchmarks such as Huang et al. (2023b), which focus on spatial or visual challenges, these prompts highlight failures rooted in the encoding process itself. Since the original set includes only 30 examples, we expand it to 110 prompts using GPT-4o. See Section˜A.2 for augmentation details and the full prompt list.
To evaluate our hypothesis, we conduct a simple causal test. We generate three images for each prompt: an image from the full prompt; another from the item without any context; and one from the full prompt, but with the lexical item’s tokens replaced by their uncontextualized representations (see Fig.˜7). For each image, we assess if the image depicts the intended interpretation of the item within the prompt (see A.1.6 for more details on this evaluation process). Our findings show that in 93% of the evaluated prompts, the item is misinterpreted in the original full-prompt image but correctly interpreted in the patched versions. This suggests that the failure stems from incorrect item resolution during the textual encoding phase: the language model misrepresents the intended meaning of the item due to contextual interference, retrieving its incorrect sense. Importantly, this indicates a failure in semantic resolution that originates entirely within the text encoder, and cannot be mitigated during image generation.
Addressing such failures requires solutions during the textual encoding that target semantic disambiguation during encoding, rather than relying on downstream interventions such as Chefer et al. (2023a). Feng et al. (2023) offer one such solution, encoding items separately according to structured representations to improve attribute binding. While this represents a promising first step, future work must account for more complex interactions and employ more sophisticated, semantically-informed methods to achieve this separation.
@@
Figure 7: Verifying textual semantic leakage by replacing the contextually leaked concept representation. (1) Regular generation produces an image showing a crosswalk to the right of a bus station. (2) Generation from the prompt “standing zebra”, without any context, results in the correct interpretation of the zebra as an animal. (3) Generation using the original prompt, but with the leaked concept “standing zebra” replaced by it’s uncontextualized representation, yields a correct image.
5 Related Work
Interpretability in T2I models.
Recent work has explored how T2I models encode and align concepts. Chefer et al. (2023b) studied CLIP’s latent space, while Toker et al. (2024) interpreted representations across the textual encoding process using the diffusion model as a lens. Others examined text-to-image alignment via attention maps (Tang et al., 2023). Another direction is to apply SAEs (Cunningham et al., 2023) to interpret intermediate representations in T2I models (Cywiński and Deja, 2025). We on the other hand, focus on token-level information at the encoder’s final layer, which directly conditions generation process.
Token representation and flow in LLMs.
Studies have shown that token information in LLMs is not uniformly distributed. Kaplan et al. (2025) found that subwords fuse into word-level meaning, while Feucht et al. (2024) documented how later tokens can erase earlier ones. Attention-based flow (Vig and Belinkov, 2019; Clark et al., 2019) can be misleading (Pruthi et al., 2020; Jain and Wallace, 2019), prompting alternatives like Attention Rollout (Abnar and Zuidema, 2020). Methods like Patchscopes (Ghandeharioun et al., 2024) and logit lens (nostalgebraist, 2020) decode internal representations, but do not test if downstream components use them.
Probing and causal methods.
Probing methods (Adi et al., 2016; Liu et al., 2019; Zhang and Bowman, 2018; Brunner et al., 2019) reveal information presence but are not fully reliable, as probes can exploit spurious correlations (Belinkov, 2022). We instead use causal interventions to test whether token information is actually used during image generation.
Challenges in T2I models.
Semantic leakage—when context distorts word meaning—has been observed in both LLMs (Gonen et al., 2024) and T2I (Rassin et al., 2022; Dahary et al., 2024). Another common issue is neglect, where key items are omitted (Chefer et al., 2023a; Chang et al., 2024b). While prior work focuses on solutions during the diffusion process, we show that some of these failures often originate in the text encoder.
6 Discussion
Improving generation via token-level intervention.
Our results suggest that generation can be improved by removing non-representative tokens, which enhances prompt-image alignment and reduces noise. This intervention can be automated via a high-precision probe (see Section˜A.5), or integrated into model design, as in MrT5 (Kallini et al., 2025), which omits uninformative tokens in byte-level encoders (Xue et al., 2022). Similarly, attention-based techniques (Chefer et al., 2023a) may benefit from prioritizing representative tokens over full lexical spans.
Developing textually challenging benchmarks.
Current T2I datasets focus primarily on visual or spatial complexity (Huang et al., 2023a; Ghosh et al., 2023; Saharia et al., 2022; Yu et al., 2022b). Yet our findings show that even slight linguistic ambiguity—particularly with polysemous or compositional phrases—can cause encoding failures. This highlights the need for evaluation benchmarks that probe textual difficulty more directly, which may in turn drive improvements in encoder design.
Generalization across T2I architectures.
We repeat our main experiments on SDXL-Turbo and Sana (see A.3.2 and A.3.3), which use unidirectional text encoders, and on FLUX-dev (A.3.1), a compute-intensive variant of FLUX. We observe similar phenomena across models, indicating our findings are not specific to a particular T2I design.
7 Conclusion
Our investigation of text encoding in T2I models revealed two key properties: (1) lexical meaning tends to concentrate in one or two representative tokens rather than distributing evenly, and (2) contextual information can leak between lexical items, resulting in misinterpretation. Our analysis further indicates focused evaluation of the text encoder and token-level interventions are promising avenues towards more semantically aligned T2I generation.
Limitations
Evaluating token-level representations remains challenging. While we rely on strong vision-language models as judges and validate key findings through human evaluation, these are still approximations of true semantic alignment. Our prompt set focuses on object-centric, syntactically simple cases, which may limit generalization to prompts involving misspellings, rare words, or abstract concepts. Further work is needed to explore how information flow behaves under more linguistically complex conditions.
@@ A.1 Technical Details
A.1.1 Lexical Item Classification
@@
A.1.2 Redundant Token Classification
We propose a probing classifier to predict whether a token is redundant (i.e., non-representative of its lexical item) using solely its encoded representation, without the need to generate an image. For this purpose, we extracted the 6,966 tokens corresponding to 4,864 unique lexical items from our dataset. Each token was annotated with a binary label indicating whether it represents the lexical item by the VLM (see A.1.3 for more details).
We split the data into training and validation sets using an 80-20 ratio. A $k$-nearest neighbors (k-NN) classifier with $k=5$ and Euclidean distance as the similarity measure was then applied to predict token redundancy directly from the encoded representations. The results on the evaluation set are presented in Table˜3. The high precision indicates that, in practical settings, one can remove tokens predicted as redundant with a high degree of certainty that representative tokens will not be inadvertently discarded.
@@ A.1.2 Redundant Token Classification
A.1.3 Evaluation Visual Generations.
@@
"In Yes, No and maybe. Does every image match one of those descriptions: (description string)? Answer Yes if all images match or relate to at least one description, Maybe if only some match, otherwise No."
@@
A.1.4 Evaluating Generated Textual Descriptions.
We employ GPT-4o (OpenAI et al., 2024) to evaluate the textual interpretations produced by Patchscopes. We use the following prompt:
@@
A.1.5 Evaluating Relations Between Items.
We enhance our leakage validation by distinguishing between cases where two lexical items exhibiting semantic leakage are perceptually bound together—for example, “old” and “man” in the prompt “a portrait of an old man”—and cases where they are not as “cone hat” and “eating” in the prompt “A person searing a cone hat is eating” (see Fig.˜6). To achieve this, we use a large language model (LLM) as a judge. Specifically, we use GPT-4o and employ the following prompt:
@@
A.1.6 Intended Item Evaluation
Foe each lexical-item, we manualy create two interpretation - one in the intended interpretation in the prompt, and another is a possible wrong interpretation of the word in other contexts. For example, given the prompt “A standing zebra to the right of a city bus station”, the corrent interpretations would be “the animal zebra”, while the incorrect interpretations would be “A zebra crossing”. We then ask a VLM to evaluate the generated image, and asses if the first or the latter interpretations exists in the images.
To evaluate the model’s capacity for contextual disambiguation, we manually define two interpretations for each lexical item in each prompt. The first is the intended semantic meaning derived from the prompt’s context, and the second is a alternative interpretation, that is wrong in this context. For instance, in the prompt “A standing zebra to the right of a city bus station”, the intended meaning of “zebra is the animal”, whereas the wrong interpretation is a “zebra crossing”. Subsequently, a VLM analyzes the generated image to determine if it depicts the intended interpretation or the wrong interpretation as we define it.
A.1.7 Resources
Our computational experiments involved inference with four distinct text-to-image models: Flux-dev, Flux-schnell, sdxl-turbo, and Sana. The parameter sizes for these models are approximately 12 billion for Flux-dev and Flux-schnell, 3.1 billion for sdxl-turbo, and a range of 0.6 to 4.8 billion for the Sana models, with our experiments utilizing a 1.6 billion parameter version. The total computational budget for these experiments is estimated to be approximately 480 GPU hours. The computing infrastructure consisted of a cluster of eight NVIDIA A100 GPUs. This configuration provided the necessary computational power for the large number of inference tasks performed.
A.1.8 Use of AI Assistants.
We utilized AI assistants to support this research. For coding the experiments, we used Microsoft’s Copilot and Anthropic’s Claude 3; all generated code was manually reviewed and validated by us to ensure it aligned with our requirements. For the paper, Google’s Gemini models (Pro and Flash) were used to improve the writing and clarity. We have carefully reviewed all content to ensure it accurately reflects our intentions.
@@ A.2 Data
DrawBench (Saharia et al., 2022): We include all categories except for “misspelling”, “rare words”, and “text”. Overall we extract 134 prompts from DrawBench.
Parti Prompts (Yu et al., 2022a): We include all categories except “Style & Format”, “Writing & Symbols”, and “Arts”. Overall we extract 923 prompts from Parti Prompts. The dataset is released under apache-2.0 license.
In total, we obtain 1,056 prompts.
Extended dataset of leakage prompts.
Our augmentation process incorporates two components. First, we generate variations of existing prompts from Rassin et al. (2022) (e.g., modifying ’a gentleman with a bow in the forest’ to ’a man wearing a bow in the jungle’). Second, we introduce novel prompts with potential semantic leakage. For these prompts, we applied a one-lexical item change test by generating an image from a similar prompt that substitutes the affected or leaked item with an alternative term (e.g., replacing ’bishops’ with ’cardinals’ or ’checkers’ in ’chess in ’2 bishops playing chess’). This test ensures that minimal lexical modifications do not alter the intended semantic meaning while producing a different image due to semantic leakage from another item in the prompt (see the first two columns in Fig.˜8 for few visual examples). Together, these methods enrich the dataset and provide a robust framework for analyzing semantic leakage. The full list of prompts is available in an anonymous Git repository. [^3] https://anonymous.4open.science/r/ TokenRole-3E19
@@
We split the data into training and validation sets using an 80-20 ratio. A $k$-nearest neighbors (k-NN) classifier with $k=5$ and Euclidean distance as the similarity measure was then applied to predict token redundancy directly from the encoded representations. The results on the evaluation set are presented in Table˜3. The high precision indicates that, in practical settings, one can remove tokens predicted as redundant with a high degree of certainty that representative tokens will not be inadvertently discarded.
A.3 Additional models
A.3.1 FLUX-Dev
In addition to our primary experiments with FLUX, we repeated all analyses using the Flux-dev variant. The redundant versus representative token experiments yielded similar trends, with 55% of tokens identified as representative and 45% as non-representative—values closely matching those observed with FLUX. Likewise, our inter-item flow experiments confirmed that information flow occurred in 11% of cases (and 3.1% miss intended leakage), reinforcing the overall patterns reported in the main text. Notably, while the aggregate trends are consistent across models, the specific lexical items resolved can differ between FLUX-schnell and Flux-dev, indicating potential a slightly different inner-lexicon (Kaplan et al., 2025). These findings underscore the robustness of our approach while highlighting model-dependent nuances in token representation and information flow dynamics.
A.3.2 SDXL-Turbo